Many companies use a quality control technique called acceptable sampling to monitor incoming shipments of parts,...

  1. Many companies use a quality control technique called acceptable sampling to monitor incoming shipments of parts, raw materials, and so on. In the electronics industry, components are commonly shipped from suppliers in large lots. Reynolds Electronics accepts a lot from a particular supplier if the defective components in the lot do not exceed 1%. Suppose a random sample of five items from a recent shipment is tested.
  1. Assume that 1 % of the shipment is defective. Compute the probability that no items in the sample are defective.
  2. Assume that 1 % of the shipment is defective. Compute the probability that exactly one item in the sample is defective.
  3. What is the probability of observing one or more defective items in the sample if 1% of the shipment is defective?
Answer #1

Let X be the random variable denoting the number of defective components in a lot of 5 items. X ~ Bin(5, 0.01)

a. The probability that no items in the sample are defective = P(X = 0) = = 0.9510

b. The probability that exactly one item in the sample is defective = P(X = 1) = = 0.0480

c. The probability of observing one or more defective items in the sample = P(X >=1) = 1 - P(X = 0) = 1 - 0.9510 = 0.0490